# NVINS: Robust Visual Inertial Navigation Fused with NeRF-augmented Camera Pose Regressor and Uncertainty Quantification

### v1

@@ V-B Uncertainty-based Outlier Rejection
As shown in Fig. 2, areas with little or no coverage in the original, unaugmented dataset $\mathcal{D}_{\text{NeRF}}$ cannot be reconstructed accurately. We find that the pose and uncertainty $\hat{\Sigma}_{n,(\cdot)}$ predicted by our BNN can be unreliable in such cases. Thus, we propose two simple and effective rejection heuristic to detect and discard pose regressor outputs which are deemed unreliable:
@@ VI Experiments
Experiments are conducted to validate (a) improved pose accuracy achieved through increased data with NeRF and uncertainty quantification, (b) consistent pose estimation without the need for loop closure, and (c) computational efficiency of the proposed VIO framework. All experiments are conducted within the photorealistic FlightGoggles simulation environment [18].
We compare the accuracy of the camera pose regressors trained with and without uncertainty prediction. Additionally, we assess the computational efficiency of the designed pose regressors on both a desktop and a Jetson AGX Xavier onboard machine (Sec. VI-A). The accuracy of our proposed VIO framework is then measured under different settings regarding uncertainties (Sec. VI-B).
We collected 1,297 posed 480 px $\times$ 480 px RGB-D images within a 15m×$\times$24m×$\times$4m region of the “Stata Center” model available in the Flightgoggles simulation environment [18]. This dataset is used to train the depth-nerfacto model [29], which incorporates the multiresolution hash encoding from Instant-NGP [13] for improved performance, and can make use of depth information, as RGB-D images are available in our case.
Upon completing the training of the NeRF model, 50,000 pairs of poses and images were generated by the trained NeRF model and utilized to train the camera pose regressors. The orientation of the sampled poses is constrained in order to avoid extreme attitudes exceeding pitch or roll angles of approximately 30 degrees. The camera pose regressor architecture includes 4 CoordConv layers [31] followed by 2 (resp. 3) fully connected layers for the pose (resp. uncertainty) output. We constructed a relatively shallow network compared to other pose regressors [20, 25, 37] to enable real-time onboard inference. Training of the pose regressor networks was conducted using the Adam optimizer with a learning rate of $10^{-3}$ for 100 epochs. The weighting factor $\kappa$ (see (8)) was set to 1. The weight decay factor for all networks is set to $\lambda=10^{-5}$.
@@ VI-A Camera Pose Regressor with Uncertainty
We trained pose regressors with two distinct loss functions: $\mathcal{L}_{total}(\mathbf{w})$ as defined in (2) and $\mathcal{L}_{2}(\mathbf{w})$ which is identical to $\mathcal{L}_{total}(\mathbf{w})$ except that the position loss and log-determinant of $\mathcal{L}(\mathbf{w})$ is replaced by the L2 loss. For each loss function, four camera pose regressors were trained and named as follows: no-nerf, vanilla, dropout, and ensemble. ‘No-nerf’ and ‘vanilla’ are neural networks trained with the dataset used to train the NeRF model and the dataset rendered by NeRF, respectively. ‘Dropout’ and ‘ensemble’ are neural networks designed to quantify the uncertainties of the outputs, utilizing MC dropout and deep ensemble, respectively. For the dropout implementation, the dropout rate was uniformly set at $p=0.05$ for all layers. ‘Ensemble’ has 8 independent neural networks are trained, each with their parameters initialized according to a Gaussian distribution with a standard deviation of $\alpha=0.02$. To ensure a fair comparison, ‘dropout’ had 8 forward passes for inference.
@@ VI-B Integration VIO with Camera Pose Regressor
Setup. Each trained pose regressor provided the VIO framework with a sequence of estimated camera poses and their uncertainties, $\{\bar{T}_{n},\hat{\Sigma}_{n}\}_{n\in\mathcal{N}}$ used to compute the sum of residual norms of the residual functions (17, 18). We employed and ‘ensemble’ regressors trained with $\mathcal{L}_{total}(\mathbf{w})$.
@@
Figure 5: Absolute Trajectory Error (ATE) for VIO frameworks with and without camera pose regressors over the travel distance of ’Trajectory 1’. ‘d’ denotes ‘dropout’, ‘e’ denotes ‘ensemble’, ‘+u’ indicates uncertainty prediction is applied, ‘+r’ indicates outlier rejection is applied, and ‘-’ signifies the absence of the respective feature.
@@
Rotational errors are lower when ‘rejection’ configurations are applied, compared to ‘no rejection’, which is noteworthy as it suggests that position-based rejection also enhances rotational performance. This is because there exists the correlation between positional and rotational uncertainties, as indicated in [21]. Furthermore, ‘rejection’ scenarios exhibit rotational accuracies that are relatively comparable to the ‘Only VIO’ scenario, while other configurations tend to underperform relative to ‘Only VIO’ with respect to rotation. However, ‘estimated’ scenarios do not consistently outperform the ‘constant’ scenarios in terms of rotational accuracy.

### v2

@@ V-B Uncertainty-based Outlier Rejection
As shown in Fig. 2, areas with little or no coverage in the original, unaugmented dataset $\mathcal{D}_{\text{NeRF}}$ cannot be reconstructed accurately. We find that the pose and uncertainty $\hat{\Sigma}_{n,(\cdot)}$ predicted by our BNN can be unreliable in such cases. Thus, we propose two simple and effective rejection heuristics to detect and discard pose regressor outputs that are deemed unreliable:
@@ VI Experiments
Experiments are conducted to validate (a) improved pose accuracy achieved through increased data with NeRF and uncertainty quantification, (b) consistent pose estimation without the need for loop closure, and (c) computational efficiency of the proposed VIO framework. All experiments are conducted within the FlightGoggles simulation [18].
We compare the accuracy of the camera pose regressors trained with and without uncertainty prediction. We also assess the computational efficiency of the designed pose regressors on both a desktop and a Jetson AGX Xavier onboard machine (Sec. VI-A). The accuracy of our proposed VIO framework is then measured under different settings regarding uncertainties (Sec. VI-B).
We collected 1,297 posed 480 px×$\times$480 px RGB-D images within a 15m×$\times$24m×$\times$4m region of the “Stata Center” model available in the Flightgoggles simulation [18]. This dataset is used to train the depth-nerfacto model [29], which incorporates the multiresolution hash encoding from Instant-NGP [13] for improved performance and can make use of depth information.
Upon completing the training of the NeRF model, 50,000 pairs of poses and images were generated by the trained NeRF model and utilized to train the camera pose regressors. The orientations of the sampled poses are constrained in order to avoid extreme attitudes exceeding pitch or roll angles of approximately 30 degrees. The camera pose regressor architecture includes 4 CoordConv layers [32] followed by 2 (resp. 3) fully connected layers for the pose (resp. uncertainty) output. We constructed a relatively shallow network compared to other pose regressors [20, 25, 38] to enable real-time onboard inference. Training of the pose regressor networks was conducted using the Adam optimizer with a learning rate of $10^{-3}$ for 100 epochs. The weighting factor $\kappa$ (see (7)) was set to 1. The weight decay factor for all networks is set to $\lambda=10^{-5}$. These parameters were selected through trial and error to maximize the performance of the networks.
@@ VI-A Camera Pose Regressor with Uncertainty
We trained pose regressors with two distinct loss functions: $\mathcal{L}_{total}(\mathbf{w})$ as defined in (2) and $\mathcal{L}_{2}(\mathbf{w})$ which is identical to $\mathcal{L}_{total}(\mathbf{w})$ except that the positional covariance $S_{n,p}^{a}$ of $\mathcal{L}_{2}(\mathbf{w})$ is considered constant. For each loss function, four camera pose regressors were trained and named as follows: no-nerf, vanilla, dropout, and ensemble. ‘No-nerf’ and ‘vanilla’ are neural networks trained with the dataset used to train the NeRF model and the dataset rendered by NeRF, respectively. ‘Dropout’ and ‘ensemble’ are neural networks designed to quantify the uncertainties of the outputs, utilizing MC dropout and deep ensemble, respectively. For the dropout implementation, the dropout rate was uniformly set at $p=0.05$ for all layers. ‘Ensemble’ has 8 independent neural networks trained, each with their parameters initialized according to a Gaussian distribution with a standard deviation of $\alpha=0.02$. To ensure a fair comparison, ‘dropout’ had 8 forward passes for inference.
@@ VI-B Integration VIO with Camera Pose Regressor
Setup. Each trained pose regressor provided the VIO framework with a sequence of estimated camera poses and their uncertainties, $\{\bar{T}_{n},\hat{\Sigma}_{n}\}_{n\in\mathcal{N}}$ used to compute the sum of residual norms of the residual functions (15, 16). We employed ’dropout’ and ‘ensemble’ regressors trained with $\mathcal{L}_{total}(\mathbf{w})$.
@@
Figure 5: Absolute Trajectory Error (ATE) for VIO with and without camera pose regressors over the travel distance of ’Trajectory 1’. ‘d’ (resp. ’e’) denotes ‘dropout’ (resp. ‘ensemble’), ‘+u’ (resp. ‘+r’) indicates uncertainty prediction (resp. outlier rejection) is applied, and ‘-’ signifies the absence of the respective feature.
@@
Rotational errors are lower when ‘rejection’ configurations are applied, compared to ‘no rejection’, which is noteworthy as it suggests that position-based rejection also enhances rotational performance. This is because there exists a correlation between positional and rotational uncertainties, as indicated in [21]. Furthermore, ‘rejection’ scenarios exhibit rotational accuracies that are relatively comparable to the ‘Only VIO’ scenario, while other configurations tend to underperform relative to ‘Only VIO’ with respect to rotation. However, ‘estimated’ scenarios do not consistently outperform the ‘constant’ scenarios in terms of rotational accuracy.